# Gaussian Process Upper Confidence Bound Achieves Nearly-Optimal Regret in Noise-Free Gaussian Process Bandits

### v1

@@ 2 Preliminaries
3 Refined Regret Upper Bound of Noise-Free GP-UCB
The following theorem describes our main results, which show the nearly optimal regret upper bound for GP-UCB.
Theorem 1 (Refined cumulative regret upper bound of GP-UCB).
Fix any input domain $\mathcal{X}\subset[-L,L]^{d}$ for some $L>0$. Suppose $B$, $L$, $d$, $\ell$, and $\nu$ are fixed constants. Then, when running Algorithm 1 under Assumption 1, the following two statements hold for any $T\in\mathbb{N}_{+}$:
  • If $k=k_{\mathrm{SE}}$, $R_{T}=O(1)$.
  • If $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$,
    $R_{T}=\begin{cases}\widetilde{O}\left(T^{\frac{d-\nu}{d}}\right)~{}~{}&\mathrm{if}~{}~{}d>\nu,\\ O\left((\ln T)^{2}\right)~{}~{}&\mathrm{if}~{}~{}d=\nu,\\ O\left(1\right)~{}~{}&\mathrm{if}~{}~{}d<\nu.\end{cases}$ (8)
Theorem 2 (Refined simple regret upper bound of GP-UCB).
Fix any input domain $\mathcal{X}\subset[-L,L]^{d}$ for some $L>0$. Suppose $B$, $L$, $d$, $\ell$, and $\nu$ are fixed constants. Then, when running Algorithm 1 under Assumption 1, the following two statements hold with $\widehat{\bm{x}}_{T}\in\operatornamewithlimits{argmax}_{\bm{x}\in\{\bm{x}_{1},\ldots,\bm{x}_{T}\}}f(\bm{x})$:
  • If $k=k_{\mathrm{SE}}$, $r_{T}=O\left(\sqrt{T}\exp\left(-\frac{1}{2}\widetilde{C}_{\mathrm{SE}}T^{\frac{1}{d+1}}\right)\right)$, where $\widetilde{C}_{\mathrm{SE}}=(6C_{\mathrm{SE}})^{-1/(d+1)}$.
  • If $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$, $r_{T}=O\left(T^{-\frac{\nu}{d}}(\ln T)^{\frac{\nu}{d}}\right)$.
Here, $C_{\mathrm{SE}}>0$ is the implied constant of the upper bound of MIG for $k=k_{\mathrm{SE}}$, which is formally defined in Lemma 3.
@@ 3 Refined Regret Upper Bound of Noise-Free GP-UCB
Our key technical results are the new analysis of the cumulative posterior standard deviation $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ and its minimum $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$, which plays an important role in the theoretical analysis of GP bandits. Indeed, following the standard analysis of GP-UCB, we have the following upper bounds of regrets by combining the UCB-selection rule with the existing noise-free confidence bound (e.g., Lemma 11 in (Lyu et al., 2019) or Proposition 1 in (Vakili et al., 2021a)):
@@
From the above inequalities, we observe that the tighter upper bounds of $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ and $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ directly result in the tighter regret upper bounds of GP-UCB. The following lemma is our main technical contribution, which gives the new upper bounds of $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ and $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$.
@@
Fix any $L>0$, input domain $\mathcal{X}\subset[-L,L]^{d}$, and kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ that satisfies $k(\bm{x},\bm{x})\leq 1$ for all $\bm{x}\in\mathcal{X}$. Furthermore, let $C_{\mathrm{SE}}$, $C_{\mathrm{Mat}}$, $\underline{\lambda}_{\mathrm{SE}}$, $\underline{\lambda}_{\mathrm{Mat}}>0$, $\underline{T}_{\mathrm{SE}}$, $\underline{T}_{\mathrm{Mat}}\geq 2$ be the constants [^2] The existence of these constants are guaranteed by the upper bound of MIG (Vakili et al., 2021b), which shows $\gamma_{T}(\lambda^{2})=O(\ln^{d+1}(T/\lambda^{2}))$ and $\gamma_{T}(\lambda^{2})=O((T/\lambda^{2})^{\frac{d}{2\nu+d}}\ln^{\frac{2\nu}{2\nu+d}}(T/\lambda^{2}))$ (as $T\rightarrow\infty,\lambda\rightarrow 0$) under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. Note that these constants do not depend on $T$, but may depend on $L$, $d$, $\ell$, and $\nu$. that satisfies $\forall\lambda\in(0,\underline{\lambda}_{\mathrm{SE}}],\forall t\geq\underline{T}_{\mathrm{SE}},\gamma_{t}(\lambda^{2})\leq C_{\mathrm{SE}}(\ln(t/\lambda^{2}))^{d+1}$ and $\forall\lambda\in(0,\underline{\lambda}_{\mathrm{Mat}}],\forall t\geq\underline{T}_{\mathrm{Mat}},\gamma_{t}(\lambda^{2})\leq C_{\mathrm{Mat}}(t/\lambda^{2})^{\frac{d}{2\nu+d}}(\ln(t/\lambda^{2}))^{\frac{2\nu}{2\nu+d}}$ for $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. Then, the following statements hold for any $T\in\mathbb{N}_{+}$ and any input sequence $\bm{x}_{1},\ldots,\bm{x}_{T}\in\mathcal{X}$:

### v2

@@ 2 Preliminaries
3 Refined Regret Upper Bound for Noise-Free GP-UCB
The following theorem describes our main results, which show the nearly-optimal regret upper bound for GP-UCB.
Theorem 1 (Refined cumulative regret upper bound for GP-UCB).
Fix any compact input domain $\mathcal{X}\subset\mathbb{R}^{d}$. Suppose $B$, $d$, $\ell$, and $\nu$ are fixed constants. Then, when running Algorithm 1 under Assumption 1, the following two statements hold for any $T\in\mathbb{N}_{+}$:
  • If $k=k_{\mathrm{SE}}$, the regret $R_{T}$ satisfies $R_{T}=O(1)$.
  • If $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$, the regret $R_{T}$ satisfies
    $R_{T}=\begin{cases}\widetilde{O}\left(T^{\frac{d-\nu}{d}}\right)~~&\mathrm{if}~~d>\nu,\\ O(\ln^{2}T)~~&\mathrm{if}~~d=\nu,\\ O\left(1\right)~~&\mathrm{if}~~d<\nu.\end{cases}$ (8)
The implied constants may depend on $B$, $d$, $\ell$, $\nu$, and the diameter of $\mathcal{X}$.
Theorem 2 (Refined simple regret upper bound for GP-UCB).
Fix any compact input domain $\mathcal{X}\subset\mathbb{R}^{d}$. Suppose $B$, $L$, $d$, $\ell$, and $\nu$ are fixed constants. Then, when running Algorithm 1 under Assumption 1, the following two statements hold by setting the estimated maximizer $\widehat{\bm{x}}_{T}$ as $\widehat{\bm{x}}_{T}\in\operatornamewithlimits{argmax}_{\bm{x}\in\{\bm{x}_{1},\ldots,\bm{x}_{T}\}}f(\bm{x})$:
  • If $k=k_{\mathrm{SE}}$, $r_{T}=O\left(\sqrt{T}\exp\left(-\frac{1}{2}CT^{\frac{1}{d+1}}\right)\right)$.
  • If $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$, $r_{T}=\widetilde{O}\left(T^{-\frac{\nu}{d}}\right)$.
The implied constants depend on $B$, $d$, $\ell$, $\nu$, and the diameter of $\mathcal{X}$. Furthermore, the constant $C>0$ depends on $d$, $\ell$, $\nu$, and the diameter of $\mathcal{X}$ [^3] As described in Section 3.1, the constant $C$ arises from the implied constants in the upper bound of MIG (vakili2021information), which depends on $d$, $\ell$, $\nu$, and the diameter of $\mathcal{X}$  [^4] Our results (Theorems 1 and 2) heavily rely on the upper bound of the MIG provided in (vakili2021information), which relies on the uniform boundness assumption of the eigenfunctions of the kernel. The validity of the uniform boundness assumption is doubted by janz2022sequential under a general compact input domain $\mathcal{X}$. Although our results are based on the upper bound of MIG in (vakili2021information), our proof strategy is also applicable for deriving nearly-optimal regrets based on the recent analysis of MIG (iwazaki2025improved) without the uniform boundness assumption. Specifically, then, the orders of the resulting regrets are the same as those in Theorems 1 and 2 for $k_{\mathrm{SE}}$ and $k_{\mathrm{Mat\acute{e}rn}}$ with $d<\nu$. As for the case under $k=k_{\mathrm{Mat\acute{e}rn}}$ with $d\geq\nu$, we can also obtain $R_{T}=\widetilde{O}(T^{(d-\nu)/d})$ and $r_{T}=\widetilde{O}(T^{-\nu/d})$, while they suffer from additional polylogarithmic terms. .
@@ 3 Refined Regret Upper Bound for Noise-Free GP-UCB
Our key technical results are the new analysis of the cumulative posterior standard deviation $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ and its minimum $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$, which plays an important role in the theoretical analysis of GP bandits. Indeed, following the standard analysis of GP-UCB, we have the following upper bounds of regrets by combining the UCB-selection rule with the existing noise-free confidence bound (e.g., Lemma 11 in (lyu2019efficient) or Proposition 1 in (vakili2021optimal)):
@@
From the above inequalities, we observe that the tighter upper bounds of $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ and $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ directly yield the tighter regret upper bounds of GP-UCB. Lemma 3 below is our main technical contribution, which gives the refined upper bounds of $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ and $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$.
@@
Fix any compact input domain $\mathcal{X}\subset\mathbb{R}^{d}$, and kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ that satisfies $k(\bm{x},\bm{x})\leq 1$ for all $\bm{x}\in\mathcal{X}$. Then, the following statements hold for any $T\in\mathbb{N}_{+}$ and any input sequence $\bm{x}_{1},\ldots,\bm{x}_{T}\in\mathcal{X}$:
  • For $k=k_{\mathrm{SE}}$, we have
    $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=O\left(\sqrt{T}\exp\left(-\frac{1}{2}CT^{\frac{1}{d+1}}\right)\right)~~\mathrm{and}~~\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=O(1).$ (11)
  • For $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$, we have
    $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ $=O\left(T^{-\frac{\nu}{d}}\ln^{\frac{\nu}{d}}T\right),$ (12)
    $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ $=\begin{cases}O\left(T^{\frac{d-\nu}{d}}\ln^{\frac{\nu}{d}}T\right)~~&\mathrm{if}~~d>\nu,\\ O\left(\ln^{2}T\right)~~&\mathrm{if}~~d=\nu,\\ O(1)~~&\mathrm{if}~~d<\nu.\end{cases}$ (13)
The constant $C>0$ and the implied constants depend on $d$, $\ell$, $\nu$, and the diameter of $\mathcal{X}$.
The full proof of Lemma 3 is given in Appendix A.2. We will also provide the proof sketch in the next section. Combining the above equations with Eqs. (9) and (10), we obtain the statements in Theorems 1 and 2.
Relation to the existing research in noise-free setting.
Lemma 3 resolves the open problem raised by the existing noise-free setting literature (liregret; vakili2022open). First, vakili2022open conjectured that the quantity $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ under $k=k_{\mathrm{Mat\acute{e}rn}}$ can attain the following upper bound:
$\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ $=\begin{cases}O\left(T^{\frac{d-\nu}{d}}\right)~~&\mathrm{if}~~d>\nu,\\ O\left(\ln T\right)~~&\mathrm{if}~~d=\nu,\\ O(1)~~&\mathrm{if}~~d<\nu.\end{cases}$ (14)
Although the above conjecture is partially validated under a specific non-adaptive algorithm (liregret; iwazaki2025improvedregretanalysisgaussian; salgiarandom), the correctness of this conjecture under a general algorithm has been an open problem (liregret). Lemma 3 answers this open problem with Eq. (13), which matches conjectured upper bound up to polylogarithmic factors.
Generality of Lemma 3.
We would like to highlight that Lemma 3 always holds for any input sequence, in contrast to the existing algorithm-specific upper bounds (iwazaki2025improvedregretanalysisgaussian; salgiarandom). Since the existing noisy GP bandits theory often leverage the upper bound of $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ or $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ from (srinivas10gaussian), we expect that many existing theoretical results in the noisy setting can be extended to the corresponding noise-free setting by directly replacing the existing noisy upper bounds of (srinivas10gaussian) with Lemma 3. For example, the analysis for GP-Thompson sampling (GP-TS) (chowdhury2017kernelized), GP-UCB and GP-TS under Bayesian setting (Russo2014-learning; srinivas10gaussian), contextual setting (krause2011contextual), GP-based level-set estimation (gotovos2013active), multi-objective setting (JMLR:v17:15-047), robust formulation (bogunovic2018adversarially), and so on.
Extension to other kernel functions.
Lemma 3 is limited to the SE and Matérn kernels. However, our proof strategy in Section 3.1 can be applied to any kernel function if we know the joint dependence of $T$ and the noise-variance parameter $\lambda^{2}$ in MIG. For example, we can apply our proof strategy for the neural tangent kernel (NTK) (jacot2018neural) by leveraging the existing upper bound of MIG under NTK (iwazaki2024no; kassraie2022neural; kassraie2022graph; vakili2021uniform). We expect that such an extension will benefit the theoretical guarantees of neural network-based bandit algorithms under a noise-free setting and will be one of the interesting research directions.
3.1 Proof Sketch of Lemma 3
In this subsection, we describe the proof sketch of Lemma 3, while we give its full proof in Appendix A.2. Below, to prove Lemma 3, we consider the more general lemmas, which bridge the MIG to noise-free posterior standard deviations.
These lemmas hold for any kernel k satisfying $k(\bm{x},\bm{x})\leq 1$, which includes most standard kernels after normalization. Roughly speaking, the above lemmas suggest that $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\lesssim\lambda_{T}$ and $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\lesssim\sum_{t=1}^{T}\lambda_{t}$ holds as far as the corresponding MIG $\gamma_{t}(\lambda_{t}^{2})$ does not increase super-linearly. Note that the MIG $\gamma_{t}(\lambda_{t}^{2})$ monotonically increases as $\lambda_{t}^{2}$ decreases, which implies that the tightest upper bound is obtained by setting $\lambda_{t}^{2}$ as $\gamma_{t}(\lambda_{t}^{2})=(t-1)/3$. By relying on the existing upper bound of MIG (vakili2021information), we can confirm that the condition $\forall t\geq\overline{T},\gamma_{t}(\lambda_{t}^{2})\leq(t-1)/3$ of the above lemmas holds with $\lambda_{t}^{2}=O(t\exp(-Ct^{\frac{1}{d+1}}))$ and $\lambda_{t}^{2}=O(t^{-\frac{2\nu}{d}}(\ln t)^{\frac{2\nu}{d}})$ for $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. Here, the constants $C$ and $\overline{T}$ are determined based on the implied constant of the upper bound of MIG. See Appendix A.2 for details. Lemma 3 follows from the aforementioned setting of $\lambda_{t}^{2}$, and Lemmas 4 and 5. Below, we give the proofs for Lemmas 4 and 5.
Proof of Lemma 4.
Instead of directly treating the noise-free posterior standard deviation, we study its upper bound with the posterior standard deviation of some noisy GP model. Here, let us denote $\sigma_{\lambda^{2}}^{2}(\bm{x};\mathbf{X}_{t-1})$ as the posterior variance under the noisy GP-model with the strictly positive variance parameter $\lambda^{2}>0$, which is defined as
Lemma 6 (Elliptical potential count lemma, Lemma D.9 in (flynn2024tighter) or Lemma 3.3 in (iwazaki2025improvedregretanalysisgaussian)).
Fix any $T\in\mathbb{N}_{+}$, any sequence $\bm{x}_{1},\ldots,\bm{x}_{T}\in\mathcal{X}$, and $\lambda>0$. Define $\mathcal{T}$ as $\mathcal{T}=\{t\in[T]\mid\lambda^{-1}\sigma_{\lambda^{2}}(\bm{x}_{t};\mathbf{X}_{t-1})>1\}$, where $\mathbf{X}_{t-1}=(\bm{x}_{1},\ldots,\bm{x}_{t-1})$. Then, the number of elements of $\mathcal{T}$ satisfies $|\mathcal{T}|\leq 3\gamma_{T}(\lambda^{2})$.
Proof of Lemma 5.
Overall, the proof strategy of this lemma is to repeatedly apply the proof of Lemma 4 by leveraging the monotonicity of the posterior variance against training inputs. First, if $T<\overline{T}$, Eq. (15) is clearly holds from the assumption $\forall\bm{x}\in\mathcal{X},k(\bm{x},\bm{x})\leq 1$. Hereafter, we focus on $T\geq\overline{T}$. By following the same argument of Lemma 4, we can confirm that there exists the index $\widetilde{t}_{T}\leq T$ such that $\sigma(\bm{x}_{\widetilde{t}_{T}};\mathbf{X}_{\widetilde{t}_{T}-1})\leq\sigma_{\lambda_{T}^{2}}(\bm{x}_{\widetilde{t}_{T}};\mathbf{X}_{\widetilde{t}_{T}-1})\leq\lambda_{T}$. Here, we define the new sequence $(\bm{x}_{t}^{(T-1)})_{t\in[T-1]}$ as the sequence that $\bm{x}_{\widetilde{t}}$ is eliminated from $(\bm{x}_{t})_{t\in[T]}$; namely, we set $\bm{x}_{t}^{(T-1)}=\mbox{1}\hskip-2.5pt\mbox{l}\{t<\widetilde{t}_{T}\}\bm{x}_{t}+\mbox{1}\hskip-2.5pt\mbox{l}\{t\geq\widetilde{t}_{T}\}\bm{x}_{t+1}$ for any $t\in[T-1]$. Furthermore, we define $\mathbf{X}_{t}^{(T-1)}=(\bm{x}_{1}^{(T-1)},\ldots,\bm{x}_{t}^{(T-1)})$. From this construction of $\mathbf{X}_{t}^{(T-1)}$, we can observe the following two facts:
  • For any $t<\widetilde{t}_{T}$, we have $\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=\sigma\left(\bm{x}_{t}^{(T-1)};\mathbf{X}_{t-1}^{(T-1)}\right)$, since $\bm{x}_{t}^{(T-1)}=\bm{x}_{t}$ and $\mathbf{X}_{t-1}^{(T-1)}=\mathbf{X}_{t-1}$.
  • For any $t>\widetilde{t}_{T}$, we have $\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\leq\sigma\left(\bm{x}_{t-1}^{(T-1)};\mathbf{X}_{t-2}^{(T-1)}\right)$, since $\bm{x}_{t-1}^{(T-1)}=\bm{x}_{t}$ and $\mathbf{X}_{t-2}^{(T-1)}\subset\mathbf{X}_{t-1}$ from the definition of $\mathbf{X}_{t}^{(T-1)}$.
From the above two facts, we have
Then, we observe that there exists the index $\widetilde{t}_{T-1}\leq T-1$ such that $\sigma(\bm{x}_{\widetilde{t}_{T-1}}^{(T-1)};\mathbf{X}_{\widetilde{t}_{T-1}-1}^{(T-1)})\leq\sigma_{\lambda_{T-1}^{2}}(\bm{x}_{\widetilde{t}_{T-1}}^{(T-1)};\mathbf{X}_{\widetilde{t}_{T-1}-1}^{(T-1)})\leq\lambda_{T-1}$ by the application of Lemma 6 for the new sequence $(\bm{x}_{t}^{(T-1)})$. Again, by setting $\bm{x}_{t}^{(T-2)}=\mbox{1}\hskip-2.5pt\mbox{l}\{t<\widetilde{t}_{T-1}\}\bm{x}_{t}^{(T-1)}+\mbox{1}\hskip-2.5pt\mbox{l}\{t\geq\widetilde{t}_{T-1}\}\bm{x}_{t+1}^{(T-1)}$ and $\mathbf{X}_{t}^{(T-2)}=(\bm{x}_{1}^{(T-2)},\ldots,\bm{x}_{t}^{(T-2)})$ for any $t\in[T-2]$, we have
$\sum_{t\in[T-1]}\sigma\left(\bm{x}_{t}^{(T-1)};\mathbf{X}_{t-1}^{(T-1)}\right)$ $\leq\sum_{t\in[T-1]\setminus\{\widetilde{t}_{T-1}\}}\sigma\left(\bm{x}_{t}^{(T-1)};\mathbf{X}_{t-1}^{(T-1)}\right)+\lambda_{T-1}$ (18)
$\leq\sum_{t\in[T-2]}\sigma\left(\bm{x}_{t}^{(T-2)};\mathbf{X}_{t-1}^{(T-2)}\right)+\lambda_{T-1}.$ (19)
$\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\leq\sum_{t=1}^{\overline{T}-1}\sigma\left(\bm{x}_{t}^{(\overline{T}-1)};\mathbf{X}_{t-1}^{(\overline{T}-1)}\right)+\sum_{t=\overline{T}}^{T}\lambda_{t}\leq\overline{T}-1+\sum_{t=\overline{T}}^{T}\lambda_{t},$ (20)
where the last inequality follows from $\sigma\left(\bm{x}_{t}^{(\overline{T}-1)};\mathbf{X}_{t-1}^{(\overline{T}-1)}\right)\leq k(\bm{x}_{t}^{(\overline{T}-1)},\bm{x}_{t}^{(\overline{T}-1)})\leq 1$. ∎
4 Discussion
Below, we discuss the remaining open questions in the noise-free setting.
  • Lower bound for squared exponential kernel. While our results establish near-optimality for the Matérn kernel, the optimal simple regret rate under the SE kernel remains unknown. Since the existing upper bound $O(\ln^{d+1}T)$ for MIG (vakili2021information) does not matches $O(\ln^{d/2}T)$ lower bound [^5] This is derived from the best-known cumulative regret upper bound $R_{T}=\widetilde{O}(\sqrt{T\gamma_{T}(\lambda^{2})})$ (li2022gaussian; salgia2021domain; valko2013finite), and lower bound $R_{T}=\Omega(\sqrt{T\ln^{d/2}T})$ in the noisy setting (scarlett2017lower)., we conjecture that further room for improvement also exists in the noise-free setting. Specifically, the exponent $d+1$ of the MIG is reflected in the denominator of the exponential factors in our regret Eq. (2). Therefore, we conjecture that $O(\sqrt{T}\exp(-CT^{2/d}))$ regret is the best guarantee for the simple regret in the SE kernel.
  • Constant cumulative regret in Bayesian setting. By using Lemma 3, we can prove the same cumulative regret as Eq. (1) up to a logarithmic factor in noise-free GP-UCB or GP-TS in the Bayesian setting (srinivas10gaussian; Russo2014-learning). However, the confidence width parameter in the Bayesian setting must scale as $\beta^{1/2}=O(\sqrt{\ln T})$ to construct a valid confidence bound. This leads to $O(\sqrt{\ln T})$ regret in SE and Matérn kernel with $d<\nu$ under the Bayesian setting, whereas the frequentist counterpart guarantees constant $O(1)$ regret (Theorem 1). This is counterintuitive, since, as shown in existing analyses for noisy settings (scarlett2018tight; srinivas10gaussian), Bayesian regret often achieves smaller values than the worst-case regret in the frequentist setting. An interesting direction for future work is to either design a Bayesian algorithm with constant regret or prove an $\Omega(\sqrt{\ln T})$ lower bound in the Bayesian setting.
  • GP-UCB in simple regret minimization. Our analysis shows that GP-UCB achieves nearly optimal simple and cumulative regrets under the Matérn kernel. On the other hand, several algorithms have been proposed that focus on minimizing the simple regret. One of the most well-known examples is EI, which greedily minimizes the simple regret under a Bayesian modeling assumption of GP, and has demonstrated good empirical performance in various applications (brochu2010tutorial; snoek2012practical). Based on our theoretical results, there exists no algorithm that can achieve strictly better simple regret than that of GP-UCB in the worst-case sense. Nevertheless, as far as we are aware, GP-UCB tends to exhibit inferior empirical performance compared to EI in simple regret minimization. See, Appendix B.2. It remains unclear whether this phenomena arises from constant factors or additional logarithmic terms in the theoretical regret upper bounds, or whether it reflects a more fundamental gap between worst-case analysis and empirical behavior.
5 Conclusion
This paper shows that GP-UCB achieves nearly-optimal regret by proving a new regret upper bound for noise-free GP bandits. The key theoretical component of our analysis is a tight upper bound on the posterior standard deviations of GP tailored to a noise-free setting (Lemma 3). As remarked in Section 3, Lemma 3 can be applicable beyond the analysis of GP-UCB. Specifically, we expect that many existing theoretical results for noisy GP bandit settings can be translated to the noise-free setting by replacing the existing noisy upper bound of the posterior standard deviations with Lemma 3. For this reason, we believe that our result marks an important step toward advancing the theory for noise-free GP bandit algorithms.
Appendix A Proofs for Section 3
A.1 Proof of Theorems 1 and 2
Lemma 7 (Deterministic confidence bound for noise-free setting, e.g., Corollary 3.11 in [kanagawa2018gaussian], Lemma 11 in [lyu2019efficient], or Proposition 1 in [vakili2021optimal]).
where $\mathbf{X}_{t}=(\bm{x}_{1},\ldots,\bm{x}_{t})$.
Although the remaining parts of the proofs are well-known results of GP-UCB, we provide the details for completeness. Based on the above lemma, we show Eqs. (9) and (10). Regarding $R_{T}$, we have
$R_{T}$ $=\sum_{t=1}^{T}f(\bm{x}^{\ast})-f(\bm{x}_{t})$ (22)
$=2B\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t}),$ (25)
where the first inequality follows from Lemma 7, and the second inequality follows from the UCB-selection rule for $\bm{x}_{t}$. Similarly to the case of cumulative regret, we have
$r_{T}$ $=f(\bm{x}^{\ast})-f(\widehat{\bm{x}}_{T})$ (26)
$\leq\min_{t\in[T]}f(\bm{x}^{\ast})-f(\bm{x}_{t})$ (27)
$=2B\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t}),$ (30)
where the first inequality follows from the definition of $\widehat{\bm{x}}_{T}$. Finally, the desired results are obtained by combining the above inequalities with Eqs. (11)–(13). ∎
A.2 Proof of Lemma 3
We prove the following Lemma 8, which is a detailed version of Lemma 3 including the dependence against constant factors.
Lemma 8 (Detailed version of posterior standard deviation upper bound for SE and Matérn kernel).
Fix any compact input domain $\mathcal{X}\subset\mathbb{R}^{d}$, and kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ that satisfies $k(\bm{x},\bm{x})\leq 1$ for all $\bm{x}\in\mathcal{X}$. Furthermore, let $C_{\mathrm{SE}}$, $C_{\mathrm{Mat}}$, $\underline{\lambda}_{\mathrm{SE}}$, $\underline{\lambda}_{\mathrm{Mat}}>0$, $\underline{T}_{\mathrm{SE}}$, $\underline{T}_{\mathrm{Mat}}\geq 2$ be the constants [^6] The existence of these constants are guaranteed by the upper bound of MIG [vakili2021information], which shows $\gamma_{T}(\lambda^{2})=O(\ln^{d+1}(T/\lambda^{2}))$ and $\gamma_{T}(\lambda^{2})=O((T/\lambda^{2})^{\frac{d}{2\nu+d}}\ln^{\frac{2\nu}{2\nu+d}}(T/\lambda^{2}))$ (as $T\rightarrow\infty,\lambda\rightarrow 0$) under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. Note that these constants do not depend on $T$, but may depend on $d$, $\ell$, $\nu$, and the diameter of $\mathcal{X}$. that satisfies $\forall\lambda\in(0,\underline{\lambda}_{\mathrm{SE}}],\forall t\geq\underline{T}_{\mathrm{SE}},\gamma_{t}(\lambda^{2})\leq C_{\mathrm{SE}}(\ln(t/\lambda^{2}))^{d+1}$ and $\forall\lambda\in(0,\underline{\lambda}_{\mathrm{Mat}}],\forall t\geq\underline{T}_{\mathrm{Mat}},\gamma_{t}(\lambda^{2})\leq C_{\mathrm{Mat}}(t/\lambda^{2})^{\frac{d}{2\nu+d}}(\ln(t/\lambda^{2}))^{\frac{2\nu}{2\nu+d}}$ for $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. Then, the following statements hold for any $T\in\mathbb{N}_{+}$ and any input sequence $\bm{x}_{1},\ldots,\bm{x}_{T}\in\mathcal{X}$: